OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a19 = -58, a21 = -164

OpenStudy (anonymous):

my choices are an = 896 - 53(n - 2) an = 896 - 53(n - 1) an = 896 + 53(n + 1) an = 896 - 53(n + 1)

OpenStudy (anonymous):

i dont understand how i use the numbers that were given to me

OpenStudy (midhun.madhu1987):